OpenStudy (anonymous):

what is 2(x-3)=-11 then what is x =

OpenStudy (anonymous):

distribute the 2 2x-6=-11 then add the 6 to both sides 2x-6+6=-11+6 2x=-5 divide by 2 2x/2=-5/2 x=-5/2

OpenStudy (anonymous):

thank you.